OpenStudy (anonymous):

how do i do this problem? find missing number 13-(8-2)=13-8-?

OpenStudy (anonymous):

subtract 2 from 8, then subtract that answer from 13. After that, subtract 8 from 13 and subtract that answer from what you got for the other side. Then switch the sign on both sides.

OpenStudy (anonymous):

so it will go 13-(6)=13-8-? 7=5-? 5-7=12 ?????

OpenStudy (anonymous):

or do i 13-6=13-8-? 7=5-? 2=-2 switch signs would be -2=2 ????

OpenStudy (anonymous):

7=5-? 7-5=-? 2=-? ?=-2

OpenStudy (anonymous):

ok ty